Air at 1 bar and 27°C is heated in a non flow system at constant pressure to 177°C. Heat is supplied from a constant temperature reservoir at 577 °C. The atmospheric temperature is 20 °C. Calculate the available energy.


1
Expert's answer
2021-06-28T08:53:19-0400

Heat received by the air is given by "Q = m_ac_v (T_2\u2032 \u2013 T_1\u2032)"

Say Q=600, "\implies 600 = 1× 0.81 (T_2′ – 400)"

"T_
2′ =\frac{ 600}{
3 × 0.81}
+ 400 = 646.9 K \space say \space 647 K"

Available energy with the source "(1200 – 290) × \frac{ 600}{
1200} = 455 kJ"

Change in entropy of the air "= m_ac_v
log
_e
\frac{T_2}{ T_1}

= 3 × 0.81 × log_e
\frac{647}{400}
= 1.168 kJ/K"

Unavailability of the air "= 290 \u00d7 1.168 = 338.72 kJ"

Available energy with the air "= 600 \u2013 338.72 = 261.28 kJ"